### v1

@@ 5.2 Main effects on outcomes
We next present 2SLS estimates of the effect of statutory recognition on settlement outcomes. We group outcomes into categories most relevant to local public goods and services: education, health, financial access, and community infrastructure.
Education.
Table 4: Effect of Local Urban Governance on School Provision
@@
Observations 37,140 37,096 37,145
  • Notes: Table reports two-stage least squares estimates where the instrument for ULB status is Census Town classification at the thresholds used in the definition of census towns. All specifications include controls for population, density, non-agricultural male workforce share, literacy, and caste composition, as well as district fixed effects101010state-level fixed effects yielded similar results.. Robust standard errors clustered at the district level are reported in parentheses.
  • Significance levels: *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
Table 4 presents the estimated effects of local urban governance on educational infrastructure provision. The results demonstrate substantial positive impacts across all levels of schooling. Statutory recognition leads to an increase of 13.54 additional primary schools (SE = 3.94), 7.76 additional middle schools (SE = 2.24), and 5.01 additional secondary schools (SE = 1.31). All estimates are statistically significant at the 1% level, with robust standard errors clustered at the district level. These findings reveal important patterns in how urban governance affects educational provision. The largest absolute effect occurs at the primary level, which is consistent with primary education being a foundational service that local governments prioritize. In India, primary schools are the least regulated and easiest to establish, requiring modest capital investment and fewer students to justify operation, whereas these requirements increase substantially for middle and secondary schools. The magnitude of effects decreases as we move up the educational ladder, with middle and secondary schools showing progressively smaller impacts. This gradient may reflect not only the differential regulatory barriers and capital requirements across school levels, but also differences in administrative priorities and demand patterns. Nonetheless, the fact that all three levels show significant positive effects suggests that urban governance structures facilitate comprehensive improvements in educational infrastructure at all levels, potentially leading to better educational outcomes for residents in newly urbanized areas.
Health.
Table 5: Effect of Local Urban Governance on Hospital Provision
@@
Observations 37,074 37,074
  • Notes: Two-stage least squares estimates where the endogenous regressor is ULB status, the instrument is Census Town status at the thresholds used in the official definition. All specifications include controls for population, density, literacy, caste composition, and economic activity, as well as district-level fixed effects [^11] State-level fixed effects yielded similar results.. Robust standard errors clustered at the district level are in parentheses. Significance: *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
Table 5 shows the effects of statutory recognition on health infrastructure. The estimates indicate that becoming a statutory town leads to an increase of 2.52 additional hospitals and 3.04 additional family welfare centers, both statistically significant at the 1% level. These results suggest that urban local governance significantly enhances health service availability, potentially improving access to medical care for residents.
@@
Table 6: Effect of Local Urban Governance on Financial Access
Dependent variable
Private banks Cooperative banks Agricultural credit societies
Effect of ULB 15.23*** 2.23*** 2.16***
(3.97) (0.89) (0.63)
Controls Yes Yes Yes
District FE Yes Yes Yes
Observations 37,062 37,060 37,085
  • Notes: Two-stage least squares estimates where the endogenous regressor is ULB status, the instrument is Census Town status at the thresholds used in the official definition. All specifications include controls for population, density, literacy, caste composition, and economic activity, as well as district-level fixed effects [^12] State-level fixed effects yielded similar results.. Robust standard errors clustered at the district level are in parentheses; 95% confidence intervals in brackets. Significance: *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
Table 6 presents the estimated effects of statutory recognition on financial infrastructure provision. The results show substantial positive impacts across all types of financial institutions. Statutory recognition leads to an increase of 15.39 additional private banks, 2.34 additional cooperative banks, and 2.24 additional agricultural credit societies. All estimates are statistically significant at the 1% level with robust standard errors clustered at the district level. These findings suggest that formal urban governance significantly enhances financial infrastructure access, with the largest effects observed for private banking services, potentially improving financial inclusion and economic development opportunities for residents.
Community Access.
Table 7: Effect of Local Urban Governance on Community Access
@@
  • Notes: Two-stage least squares estimates where the endogenous regressor is ULB status, the instrument is Census Town status at the thresholds used in the official definition. All specifications include controls for population, density, literacy, caste composition, and economic activity, as well as district-level fixed effects [^13] State-level fixed effects yielded similar results.. Robust standard errors clustered at the district level are in parentheses. Significance: *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
@@
Table 7 presents the estimated effects of statutory recognition on community infrastructure and access. The results show significant positive impacts on most community facilities. Statutory recognition leads to an increase of 1.07 additional public libraries (significant at the 5% level), 1.44 additional public reading rooms (significant at the 1% level), and 0.83 additional cinema halls (significant at the 10% level). Interestingly, the estimate for sports infrastructure is negative and significant at the 1% level, indicating a decrease of 5.69 facilities. This negative effect may reflect the spatial constraints and land-use tradeoffs inherent in urban development. As settlements formalize into statutory towns, open spaces previously used for sports facilities (such as playgrounds or fields) may be repurposed for higher-priority urban infrastructure like schools, hospitals, banks, or residential development. Moreover, municipal planning often consolidates multiple informal sports venues into fewer, more formal facilities, resulting in a net reduction in the count of such infrastructure even as quality or accessibility may improve. Overall, these findings suggest that formal urban governance enhances access to cultural and educational community facilities, though the transition may involve restructuring of certain types of infrastructure, particularly those requiring substantial land area.
5.3 Robustness and validity
5.3.1 Exclusion Restriction Assumption
A key assumption for the validity of our IV strategy is that the instrument (Census Town eligibility) affects outcomes only through its impact on statutory recognition, and not through other channels. To assess this exclusion restriction, we conduct robustness checks. First, for areas that remained non-statutory towns, we check the effects of census eligibility on outcomes to see if any direct effects exist absent statutory recognition. Second, we check for balance in baseline covariates around the thresholds to ensure that settlements just above and below the cutoffs are similar in observable characteristics (see Appendix).
Table 8: Reduced-Form Effects Comparison: Main Estimates vs Census Town Eligibility Effect on Outcomes for Non-Statutory Towns
Main 2SLS estimate Direct Effect
Outcome Estimate SE Estimate SE
Education
   Primary schools 13.85*** (4.03) 0.81*** (0.13)
   Middle schools 7.71*** (2.27) 0.51*** (0.10)
   Secondary schools 5.01*** (1.31) 0.30*** (0.06)
Health
   All hospitals 2.53*** (0.69) 0.16*** (0.04)
   Family welfare centers 2.99*** (0.88) 0.21*** (0.04)
Financial access
   Private banks 15.23*** (3.97) 0.92*** (0.11)
   Cooperative banks 2.23*** (0.89) 0.14*** (0.05)
   Agricultural credit societies 2.16*** (0.63) 0.12*** (0.04)
Community infrastructure
   Public libraries 1.04** (0.50) 0.10*** (0.03)
   Public reading rooms 1.44*** (0.63) 0.13*** (0.04)
   Cinema halls 0.83* (0.49) 0.07** (0.03)
   Sports infrastructure -5.69*** (1.46) -0.38*** (0.06)
  • Notes: The first two columns report the main 2SLS estimates from Tables 47, showing the effect of statutory ULB status instrumented by Census Town eligibility. The last two columns report direct effects of Census Town eligibility on outcomes for the subsample of settlements that remained non-statutory i.e. testing whether the instrument has direct effects on outcomes absent statutory recognition. All specifications include controls for population, density, literacy, caste composition, and economic activity, as well as district-level fixed effects. Robust standard errors clustered at the district level are in parentheses. Significance: *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
Table 8 presents evidence on the exclusion restriction assumption by comparing the main 2SLS estimates with reduced-form effects of Census Town eligibility among settlements that remained non-statutory. While the reduced-form estimates are statistically significant and consistently positive (or negative for sports infrastructure) across all outcomes, their magnitudes are substantially smaller than the main IV estimates. For example, the effect on primary schools is 0.81 (SE = 0.13) in the non-statutory sample compared to 13.85 (SE = 4.03) in the main specification, and the effect on private banks is 0.92 (SE = 0.11) versus 15.23 (SE = 3.97). The relative magnitudes thus provide support for the validity of our instrumental variable strategy, as the direct effects are an order of magnitude smaller than the effects mediated through actual municipalization.
@@ 6 Conclusion
This paper provides causal evidence that statutory municipalization – the extension of urban local governance to a settlement – leads to some improvements in local public goods at India’s urbanizing fringe. Using a novel multi-threshold RD design based on the Census Town criteria, we isolate the effect of formal urban status from underlying confounding factors such as population growth, density and economic changes. The results indicate that when a village becomes a town (in administrative terms), it experiences gains in infrastructure: more schools, health centers, banks, and other facilities emerge, relative to similar villages that remain under rural governance. These findings suggest that institutions and governance structures do play a role in shaping development outcomes, even at the very local level. However, the magnitudes of the effects are up for interpretation. While statistically significant, the increases in infrastructure may be considered modest or substantial, depending on the context and baseline levels of local services.
@@
The policy implications are twofold. First, there may be benefits to timely municipalization of emerging urban areas. Our paper shows that delaying the recognition of de facto urban settlements can mean missed opportunities for infrastructure investment and service delivery. If urban local bodies indeed facilitate better outcomes, then proactively converting eligible large villages into statutory towns, especially the villages that are already urban as per the Census Town criteria, and providing them fiscal support and governance framework, could help improve their overall infrastructure. Second, our results highlight the importance of state capacity and support for new ULBs. Simply declaring a settlement as a municipality is not a panacea; the improvements we document likely come from a combination of local initiatives and higher-level funding and oversight. State governments could enhance the benefits of municipalization by ensuring that newly formed ULBs receive adequate resources, technical assistance, and training to fulfill their functions. Strengthening local governance – through measures like capacity-building programs for municipal officials or incentivizing revenue generation – could further amplify the positive impacts on public goods.
@@
Our study also opens avenues for future research. One extension would be to examine longer-term and broader outcomes of municipalization. Due to data limitations, we focused on infrastructure inputs as of 2011. It would be valuable to see if these translate into improved local public goods for residents over time. For example, higher school enrollment or better health indicators in newly urbanized areas. Household survey data, such as subsequent rounds of NFHS, could be linked to our design to assess impacts on education, health, and income. Another avenue is to explore the fiscal side: how do the finances of local bodies change at the threshold? Are municipalities able to raise more revenue or attract more investment? Additionally, the heterogeneity across states hints at interesting political economy dynamics – understanding why some states are reluctant to create new towns, possibly to avoid sharing revenue or due to political patronage networks, could inform strategies to overcome those barriers. Our findings underscore that urban local governance matters: even in a context where local governments often have limited capacity, formally bringing a settlement into the urban administrative fold has certain measurable benefits for the community. As countries throughout the developing world continue to urbanize, ensuring that institutional change keeps pace with demographic change will be crucial for sustainable long-term development.
@@
Density Plots and McCrary Test
We conduct density tests around the statutory recognition thresholds to assess the validity of our RD design. Specifically, we implement the mccrary2008manipulation test for manipulation of the running variable, which examines the continuity of the density of the running variable at the cutoff.
Figure 3: Density plots of CT-definition variables and thresholds
The mccrary2008manipulation density test formally assesses whether there is evidence of manipulation of the running variable around the cutoff. The test estimates the density of the running variable separately on each side of the threshold and tests whether there is a discontinuity in the density at the cutoff. A significant discontinuity would suggest strategic sorting around the threshold, which would violate the identifying assumptions of the RD design.
Table 9 presents the results of the mccrary2008manipulation test for each of our three running variables used to define census town status. For population in 2001 (cutoff at 5,000), the test statistic is 0.754 with a p-value of 0.451. For population density in 2001 (cutoff at 400 persons per km²), the test statistic is 0.245 with a p-value of 0.807. For the share of non-agricultural male workers (cutoff at 0.75), the test statistic is 1.101-1.101 with a p-value of 0.271. In all three cases, we fail to reject the null hypothesis of no discontinuity in the density at conventional significance levels. This provides reassuring evidence that there is no systematic manipulation of these running variables around the statutory thresholds.
Table 9: McCrary Density Tests at CT Thresholds
Running Variable Cutoff T-statistic P-value
Population (2001) 5,000 0.754 0.451
Population Density (2001) 400 0.245 0.807
% Non-Agricultural Male Workers (2001) 0.75 1.101-1.101 0.271
  • Notes: This table reports results from the mccrary2008manipulation manipulation test which examines whether there is a discontinuity in the density of the running variable at the statutory threshold. The T-statistic tests the null hypothesis of no discontinuity. All tests use a triangular kernel with bandwidth selection and jackknife variance estimation. P-values greater than 0.10 indicate no evidence of manipulation at conventional significance levels.
@@
As expected in the full sample (Panel A), settlements meeting CT criteria differ substantially from those that do not—they are considerably larger, denser, more economically non-agricultural, and have higher literacy rates. However, when we restrict attention to the close-to-threshold sample (Panel B), these differences narrow substantially. While some gaps remain—particularly in population and density, which are among the threshold criteria themselves—the key observation is that settlements just above and below the thresholds are much more similar than the full sample comparison would suggest. This pattern is consistent with local randomization around the threshold and supports the identifying assumption that, conditional on being close to the CT cutoffs, assignment to CT status is quasi-random with respect to unobserved determinants of public goods provision.

### v2

@@ 4.2 Main effects on outcomes
We next present 2SLS estimates of the effect of statutory recognition on settlement outcomes. We group outcomes into categories most relevant to local public goods and services: education, health, financial access, and community infrastructure. The reduced-form estimates provide the numerator of the Wald ratio underlying our IV specification and show that crossing the CT eligibility frontier improves public goods provision across the same outcome categories [^11] The full reduced-form table is reported in Appendix Table 9..
Table 4: Effect of Local Urban Governance Status on School Provision
@@
Observations 37,143 37,099 37,148
  • Notes: Table reports two-stage least squares estimates where the endogenous regressor is statutory town status (ULB =1=1) and the excluded instrument is Census Town eligibility in 2001. The local sample restricts to settlements near the 2001 census thresholds (population $\pm 5{,}000$; density $\pm 400$; non-agricultural male share $\pm 0.20$). All specifications include controls for log population, log density, non-agricultural male workforce share, literacy rate, main worker share, and caste composition (SC and ST shares), as well as district fixed effects. Robust standard errors clustered at the district level are reported in parentheses.
  • Significance levels: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
Education.
Table 4 presents the estimated effects of statutory recognition on educational infrastructure provision. The results demonstrate substantial positive impacts across all levels of schooling. Statutory recognition leads to an increase of 13.86 additional primary schools, 7.72 additional middle schools, and 4.89 additional secondary schools. All estimates are statistically significant at the 1% level, with robust standard errors clustered at the district level. The largest absolute effect occurs at the primary level, consistent with primary education being a foundational service that local governments can expand most readily. Primary schools are also typically smaller in size and easier to open than middle or secondary schools, facilitating rapid increases in provision following statutory recognition. As we move up the educational ladder, the estimated effects become smaller, which is plausible given the higher capital requirements, larger catchment areas, and greater regulatory demands associated with middle and secondary schools. Even so, the fact that all three levels show significant positive effects suggests that ULB status facilitates broad improvements in educational infrastructure rather than a narrow shift at only one school level.
Table 5: Effect of Local Urban Governance Status on Hospital Provision
@@
Observations 37,091 37,077
  • Notes: Two-stage least squares estimates where the endogenous regressor is statutory town status (ULB =1=1) and the excluded instrument is Census Town eligibility in 2001. The local sample restricts to settlements near the 2001 census thresholds (population $\pm 5{,}000$; density $\pm 400$; non-agricultural male share $\pm 0.20$). All specifications include controls for log population, log density, non-agricultural male workforce share, literacy rate, main worker share, and caste composition (SC and ST shares), as well as district fixed effects. Robust standard errors clustered at the district level are reported in parentheses.
  • Significance levels: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
Health.
Table 5 presents the estimated effects of local urban governance status on health infrastructure provision. The results demonstrate substantial positive impacts across both health facilities. Local urban governance status leads to an increase of 2.53 additional hospitals and 3.00 additional family welfare centers. Both estimates are statistically significant at the 1% level, with robust standard errors clustered at the district level. The positive effects on both general hospitals and family welfare centers suggest that local urban governance status facilitates improvements in core health infrastructure. The somewhat larger effect on family welfare centers is noteworthy because these facilities play a central role in maternal and child health, family planning, and preventive care within India’s public health system (Dhingra and Dutta, 2011). This pattern is also plausible given that family welfare centers typically require less capital and can be expanded more quickly than full-service hospitals.
Table 6: Effect of Local Urban Governance Status on Financial Access
Dependent variable
Cooperative banks Agricultural credit societies
Effect of ULB status 4.09∗∗∗ 2.84∗∗∗
(0.97) (1.01)
Controls Yes Yes
District FE Yes Yes
Observations 37,089 37,093
  • Notes: Two-stage least squares estimates where the endogenous regressor is statutory town status (ULB =1=1) and the excluded instrument is Census Town eligibility in 2001. The local sample restricts to settlements near the 2001 census thresholds (population $\pm 5{,}000$; density $\pm 400$; non-agricultural male share $\pm 0.20$). All specifications include controls for log population, log density, non-agricultural male workforce share, literacy rate, main worker share, and caste composition (SC and ST shares), as well as district fixed effects. Robust standard errors clustered at the district level are reported in parentheses.
  • Significance levels: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
@@
Table 6 presents the estimated effects of local urban governance status on financial infrastructure provision. The results demonstrate substantial positive impacts across both types of financial institutions. Local urban governance status leads to an increase of 4.09 additional cooperative banks and 2.84 additional agricultural credit societies, both statistically significant at the 1% level with robust standard errors clustered at the district level. The positive effects across both institution types indicate that municipalization enhances multiple forms of financial intermediation serving different community needs. Cooperative banks can support broader commercial and household financial activity, while agricultural credit societies remain relevant in peri-urban settlements where agricultural livelihoods persist alongside urbanization. The expansion of financial infrastructure may therefore reflect both greater demand in newly recognized towns and the possibility that local urban governance status signals administrative capacity and development potential to formal financial institutions (Burgess and Pande, 2005).
Table 7: Effect of Local Urban Governance Status on Community Infrastructure
@@
  • Notes: Two-stage least squares estimates where the endogenous regressor is statutory town status (ULB =1=1) and the excluded instrument is Census Town eligibility in 2001. The local sample restricts to settlements near the 2001 census thresholds (population $\pm 5{,}000$; density $\pm 400$; non-agricultural male share $\pm 0.20$). All specifications include controls for log population, log density, non-agricultural male workforce share, literacy rate, main worker share, and caste composition (SC and ST shares), as well as district fixed effects. Robust standard errors clustered at the district level are reported in parentheses.
  • Significance levels: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
Community Infrastructure.
@@
Table 7 presents the estimated effects of local urban governance status on community infrastructure and access. The results show positive effects on several cultural and educational amenities. Local urban governance status leads to a statistically significant increase of 1.05 public libraries and 1.44 public reading rooms. The estimate for cinema halls is positive at 0.79 but imprecisely estimated. By contrast, sports infrastructure declines by 5.71 facilities. This negative effect is consistent with spatial constraints and land-use tradeoffs in urban development: as settlements formalize into statutory towns, open spaces previously used for sports facilities may be repurposed for schools, health facilities, banking infrastructure, roads, or residential development. Overall, these findings suggest that formal statutory recognition may enhance access to community facilities, though the transition may involve restructuring of certain types of infrastructure, particularly those requiring substantial land area.
4.3 Robustness Checks
4.3.1 Density Plots and McCrary Test for Manipulation
We conduct checks around the statutory recognition thresholds to assess the validity of our RD design. Specifically, we plot the density of the running variables and implement the McCrary (2008) test for manipulation of the running variable, which examines the continuity of the density of the running variable at the cutoff.
Figure 3: Density plots of CT-definition variables and thresholds
Figure 3 shows the density plots of the three running variables (population density, population, and share of non-agricultural male workers) around their respective cutoffs of 400 persons per km², 5,000 persons, and 0.75. The plots indicate that the density of the running variables is continuous and smooth around the cutoffs, with no clear signs of manipulation or strategic sorting.
Table 8: McCrary Density Tests at CT Thresholds
Running Variable Cutoff T-statistic P-value
Population (2001) 5,000 0.754 0.451
Population Density (2001) 400 0.245 0.807
% Non-Agricultural Male Workers (2001) 0.75 1.101-1.101 0.271
  • Notes: This table reports results from the McCrary (2008) manipulation test which examines whether there is a discontinuity in the density of the running variable at the statutory threshold. The T-statistic tests the null hypothesis of no discontinuity. All tests use a triangular kernel with bandwidth selection and jackknife variance estimation. P-values greater than 0.10 indicate no evidence of manipulation at conventional significance levels.
The McCrary (2008) density test formally assesses whether there is evidence of manipulation of the running variable around the cutoff. The test estimates the density of the running variable separately on each side of the threshold and tests whether there is a discontinuity in the density at the cutoff. A significant discontinuity would suggest strategic sorting around the threshold, which would violate the identifying assumptions of the RD design.
Table 8 presents the results of the McCrary (2008) test for each of our three running variables used to define census town status. For population in 2001 (cutoff at 5,000), the test statistic is 0.754 with a p-value of 0.451. For population density in 2001 (cutoff at 400 persons per km²), the test statistic is 0.245 with a p-value of 0.807. For the share of non-agricultural male workers (cutoff at 0.75), the test statistic is 1.101-1.101 with a p-value of 0.271. In all three cases, we fail to reject the null hypothesis of no discontinuity in the density at conventional significance levels. This provides reassuring evidence that there is no systematic manipulation of these running variables around the statutory thresholds. Remaining robustness tests can be found in the Appendix.
@@ 5 Conclusion
This paper provides evidence that local urban governance status leads to meaningful improvements in local public goods at India’s urbanizing fringe. Using a novel multi-threshold fuzzy RD design that leverages Census Town eligibility criteria, we isolate the effect of local urban governance status on public goods provision from underlying confounding factors.
Our local IV estimates show that statutory recognition increases the number of government schools, hospitals, family welfare centers, cooperative banks, and agricultural credit societies, while decreasing the number of sports facilities, consistent with the idea of reallocation of space and land constraints in formalizing settlements. These local average treatment effects indicate that ULB status materially shifts the stock of public infrastructure in settlements near the Census Town thresholds. The corresponding reduced-form estimates are directionally consistent and confirm that crossing the eligibility frontier improves public goods provision even before scaling by the first stage.
@@
The policy implications are twofold. First, there are clear benefits to timely municipalization of emerging urban areas. Our results show that delaying the recognition of urban settlements may imply missed opportunities for public infrastructure. Proactively converting eligible large villages into statutory towns and providing them with fiscal support and a governance framework could help improve their overall infrastructure. Second, our results highlight the importance of state capacity and support for new ULBs. We emphasize that simply declaring a settlement as a municipality is not a panacea. The improvements we document likely come from a combination of federal, state, and local initiatives, the ability to levy local taxes, greater funding availability, and improved local governance. State governments could enhance the benefits of municipalization by ensuring that newly formed ULBs receive adequate resources, technical assistance, and training to fulfill their functions.
@@
Our study opens avenues for future research. One extension would be to examine longer-term outcomes beyond 2011, using household survey data such as subsequent rounds of NFHS to assess impacts on education, health, and income. A clear need remains to understand why and how local urban governance status leads to improved public goods provision. Is it primarily through increased fiscal resources, improved administrative capacity, greater political representation, or a combination of these factors? We encourage future work to unpack the mechanisms through which municipalization translates into better infrastructure and services. Another avenue is to explore the fiscal channel: how do municipal finances change at the threshold, and to what extent do new ULBs rely on own-source revenue versus transfers? Understanding heterogeneity in statutory recognition across states, such as why some are reluctant to create new towns, possibly to avoid sharing revenue or due to political patronage networks, could inform policymakers seeking to overcome barriers to timely municipalization. Our findings underscore that formally bringing settlements into the urban administrative fold has measurable benefits, and that ensuring institutional change keeps pace with demographic change will be crucial for sustainable development in rapidly urbanizing economies.
@@
Reduced-Form Estimates
Table 9 reports the reduced-form effect of crossing the Census Town eligibility frontier on local public goods provision in the local sample. These estimates provide the numerator of the Wald ratio underlying the main IV results in the text. The sign pattern matches the IV estimates throughout: education, health, financial access, and selected community amenities improve, while sports infrastructure declines.
Table 9: Reduced-Form Effects of Census Town Eligibility on Public Goods Provision
Outcome Estimate SE NN
Education
   Primary schools 0.982∗∗∗ (0.212) 37,143
   Middle schools 0.572∗∗∗ (0.121) 37,099
   Secondary schools 0.346∗∗∗ (0.066) 37,148
Health
   All hospitals 0.179∗∗∗ (0.036) 37,091
   Family welfare centers 0.212∗∗∗ (0.041) 37,077
Financial access
   Cooperative banks 0.297∗∗∗ (0.044) 37,089
   Agricultural credit societies 0.202∗∗∗ (0.061) 37,093
Community infrastructure
   Public libraries 0.074∗∗ (0.030) 37,147
   Public reading rooms 0.102∗∗∗ (0.035) 37,147
   Cinema halls 0.056 (0.030) 37,146
   Sports infrastructure -0.405∗∗∗ (0.057) 37,146
Controls: Yes  District FE: Yes  SEs clustered by district
  • Notes: Reduced-form estimates from OLS regressions of each outcome on Census Town eligibility in 2001 (a binary indicator for meeting all three CT thresholds). The local sample restricts to settlements near the 2001 census thresholds (population $\pm 5{,}000$; density $\pm 400$; non-agricultural male share $\pm 0.20$). All specifications include controls for log population, log density, non-agricultural male workforce share, literacy rate, main worker share, and caste composition (SC and ST shares), as well as district fixed effects. Robust standard errors clustered at the district level are in parentheses.
  • $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$.
In magnitude, the reduced-form estimates are smaller because only a subset of CT-eligible settlements is ultimately statutorily recognized. Even so, the effects remain consistent with IV estimates: crossing the eligibility frontier raises the number of primary schools, middle schools, and secondary schools, while also increasing hospitals, family welfare centers, cooperative banks, and agricultural credit societies. The decline in sports infrastructure is also visible in reduced form, consistent with the land-reallocation interpretation discussed in the main text.
@@
As expected in the full sample (Panel A), settlements meeting CT criteria differ substantially from those that do not as they are considerably larger, denser, more economically non-agricultural, and have higher literacy rates. However, when we restrict attention to the close-to-threshold sample (Panel B), these differences narrow substantially. While some gaps remain, particularly in population and density, which are among the threshold criteria themselves, the key observation is that settlements just above and below the thresholds are much more similar than the full sample comparison would suggest. Neverthless, we control for these remaining differences in our regression models by adding these variables along with appropriate interaction terms and fixed effects.
Bandwidth Sensitivity
[…]
Table 11: Bandwidth Sensitivity: Primary School Estimates
Bandwidth NN NN(ST=1) FS coef RF est RF SE IV est IV SE
[…]
Descriptive Outcome Levels
Table 12 reports unadjusted outcome means by statutory status in the full sample and in the close-to-threshold local sample. These comparisons are purely descriptive and are not part of the identification argument. Their value is diagnostic: they show how large the raw full-sample differences are, and how much those differences compress once we restrict attention to settlements near the Census Town thresholds used in the main analysis.
Table 12: Descriptive Statistics: Outcome Variables by Statutory Status in the Global and Local Samples
Global sample Local sample
Outcome Non-ST ST Difference Non-ST ST Difference
Education
Primary schools 1.27 15.50 14.23 1.24 4.96 3.72
Middle schools 0.55 7.64 7.09 0.58 2.65 2.07
Secondary schools 0.19 4.30 4.12 0.25 1.64 1.39
Health
All hospitals 0.01 1.13 1.12 0.02 0.41 0.39
Family welfare centers 0.06 1.18 1.12 0.08 0.53 0.46
Financial access
Cooperative banks 0.06 1.90 1.83 0.09 0.97 0.88
Agricultural credit societies 0.16 2.21 2.04 0.17 0.92 0.75
Community infrastructure
Public libraries 0.13 1.02 0.89 0.11 0.43 0.31
Public reading rooms 0.15 1.03 0.88 0.13 0.45 0.32
Cinema halls 0.06 1.73 1.67 0.05 0.13 0.08
Sports infrastructure 0.42 1.06 0.64 0.46 0.49 0.03
  • Notes: Entries are unadjusted means of each outcome by statutory town status. “Difference” column is the raw difference in means between statutory and non-statutory settlements. The global sample uses the full regression sample with non-missing covariates and outcome data. The local sample uses the close-to-threshold sample from the main analysis. Outcome-specific sample sizes vary slightly because of missingness and match the corresponding samples used in Table 13. These differences are descriptive only.
The descriptive comparison reveals a sharp contrast between the two samples. In the full sample, statutory settlements have far higher levels of nearly every public good: for example, they have 15.50 primary schools versus 1.27 in non-statutory settlements, 1.90 cooperative banks versus 0.06, and 1.73 cinema halls versus 0.06. Once we move to the local sample, these gaps narrow substantially. The compression is especially stark for community amenities: the raw gap in cinema halls falls from 1.67 to 0.08, and the raw gap in sports infrastructure shrinks from 0.64 to 0.03. Note that these raw differences are not adjusted for any covariates or fixed effects, and thus should not be interpreted as causal estimates of local urban governance.
OLS versus IV Estimates
Table 13 takes the comparison one step further by moving from raw descriptive differences to adjusted regression estimates. Column (1) reports global OLS estimates from the full sample of 502,110 settlements. Column (2) restricts the OLS to the close-to-threshold local sample used in the main analysis. Column (3) reports the local 2SLS estimates from Tables 47, where statutory status is instrumented by Census Town eligibility. Reading alongside Table 12, the progression is straightforward: global raw differences are large, local raw differences are smaller, local OLS attenuates further after conditioning on observables, and the local IV estimates provide the paper’s causal parameter.
Table 13: OLS versus IV Estimates of the Effect of Statutory Recognition
(1) (2) (3)
OLS OLS 2SLS
Global Local Local
Education
     Primary schools 11.20*** 2.08*** 13.86***
(0.95) (0.49) (4.00)
     Middle schools 5.22*** 1.10*** 7.72***
(0.42) (0.22) (2.25)
     Secondary schools 3.14*** 0.92*** 4.89***
(0.24) (0.10) (1.30)
Health
     All hospitals 1.00*** 0.36*** 2.53***
(0.06) (0.06) (0.69)
     Family welfare centers 0.85*** 0.33*** 3.00***
(0.06) (0.05) (0.88)
Financial access
     Cooperative banks 1.46*** 0.64*** 4.09***
(0.10) (0.07) (0.97)
     Agricultural credit societies 1.47*** 0.47*** 2.84***
(0.26) (0.09) (1.01)
Community infrastructure
     Public libraries 0.51*** -0.01 1.05**
(0.06) (0.06) (0.50)
     Public reading rooms 0.49*** -0.08 1.44**
(0.06) (0.07) (0.63)
     Cinema halls 1.49*** -0.01 0.79
(0.11) (0.04) (0.49)
     Sports infrastructure -0.09 -0.46*** -5.71***
(0.09) (0.08) (1.48)
Controls Yes Yes Yes
District FE Yes Yes Yes
Observations 502,110 37,148 37,148
  • Notes: Columns (1) and (2) report OLS estimates of the coefficient on statutory town status ($ST_{i}$). Column (3) reports 2SLS estimates where $ST_{i}$ is instrumented by Census Town eligibility in 2001 ($Z_{i}$). All specifications include controls for log population, log density, non-agricultural male workforce share, literacy rate, main worker rate, and caste shares, as well as district fixed effects. Robust standard errors clustered at the district level are in parentheses. The “Global” sample includes all settlements with non-missing covariates. The “Local” sample restricts to settlements within $\pm 5{,}000$ of the population threshold, $\pm 400$ of the density threshold, and $\pm 0.20$ of the non-agricultural share threshold.
  • Significance: *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
Taken together, Tables 12 and 13 reveal three instructive patterns. First, both the raw means and the global OLS estimates in Column (1) show very large positive differences between statutory and non-statutory settlements for nearly all outcomes. Those differences are informative about broad urban-rural contrasts, but they are heavily contaminated by selection: larger, denser, and more economically developed settlements are both more likely to be statutory towns and more likely to have schools, hospitals, and banks, regardless of their governance status. The global comparison therefore conflates the effect of statutory recognition with pre-existing differences across places.
Second, restricting the comparison to the close-to-threshold local sample sharply compresses the differences. This is visible in the raw means above and again in the adjusted local OLS estimates in Column (2). For education outcomes, the OLS coefficient on primary schools drops from 11.20 to 2.08, and for middle schools from 5.22 to 1.10. For community infrastructure, the local OLS estimates are essentially zero — public libraries (0.01-0.01), reading rooms (0.08-0.08), and cinema halls (0.01-0.01) show no conditional association with statutory status. This attenuation is exactly what we would expect if settlements near the thresholds are much more comparable than the full sample.
Third, Column (3) shows that the local IV estimates that are interpreted as our paper’s causal estimates for settlements whose statutory recognition status was shifted by crossing the Census Town thresholds. These 2SLS coefficients are uniformly larger than the local OLS estimates, and for several outcomes they are larger than even the global OLS estimates. For example, the IV estimate for family welfare centers (3.00) is more than 3.5 times the global OLS (0.85), and the IV estimate for public libraries (1.05) is significant where the local OLS is essentially zero. This pattern is consistent with a LATE interpretation: statutory recognition has especially large effects for marginal settlements near the eligibility frontier, while naive cross-sectional comparisons either mix together incomparable places or understate the effect because of attenuation in the observed status measure.